A= x2zi + xy j +y2z k, B = yz2 +xz j + x2z k, find grad(A.B)


1
Expert's answer
2021-09-02T00:35:46-0400
"A\cdot B=x^2z(yz^2)+xy(xz)+y^2z(x^2z)"

"=x^2yz^3+x^2yz+x^2y^2z^2"

"grad(A\cdot B)=\nabla(A\cdot B)"

"=\langle\dfrac{\partial}{\partial x}(A\cdot B), \dfrac{\partial}{\partial y}(A\cdot B),\dfrac{\partial}{\partial z}(A\cdot B)\rangle"

"=(2xyz^3+2xyz+2xy^2z^2)i"

"+(x^2z^3+x^2z+2x^2yz^2)j"

"+(3x^2yz^2+x^2y+2x^2y^2z)k"